OpenStudy (anonymous):

Tomar's paycheck varies directly with the hours he works. When Tomar works 5 hours, he gets paid $60.00. How much does Tomar make in 8 hours? $75 $80 $96 $120

OpenStudy (austinl):

\(P=w\times t\) \(P=\text{Paycheck}\) \(w=\text{hourly wages}\) \(t=\text{time worked}\) We have, \(60=w(5),~\text{Solve for w.}\) Once you know w, you can plug in the value for \(t=8\)

OpenStudy (anonymous):

60 divided by five is 12 he makes 12 the hour 8 times 12 = 96

OpenStudy (austinl):

Or alternatively, \(\large{\dfrac{5}{60}=\dfrac{8}{x}}~\text{Solve for x.}\)

OpenStudy (anonymous):

thanks